Solve 11+x=-3 and check the solution
Answer by nyc_function(2741)

11 + x = -3

Subtract 11 from both sides of the equation to isolate x.

-11 + 11 + x = -3 - 11

The positive and negative 11 cancel each other out on the left side and thus isolating x.

x = -3 - 11

x = -14

THE PROOF:

Let x = -14 in the original linear equation.

11 + x = -3
11 + (-14) = -3

-3 = -3....It checks!

We can now confirm that x = -14 is correct.